40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 26 17:24:24 2017, Steve, Update, General, power outage powerOutage.png
    Reply  Thu Dec 28 17:22:02 2017, gautam, Update, General, power outage - CDS recovery 28.png
       Reply  Thu Jan 4 14:39:50 2018, gautam, Update, General, power outage - timing error 
          Reply  Fri Jan 5 21:54:28 2018, rana, Update, General, power outage - timing error huh.png
             Reply  Fri Jan 5 22:19:53 2018, gautam, Update, General, power outage - timing error 
                Reply  Sat Jan 6 18:27:37 2018, gautam, Update, General, power outage - IFO recovery EX_QPD_Quad1_Faulty.pdfTrillium_faulty.pdf
                   Reply  Mon Apr 16 22:09:53 2018, Kevin, Update, General, power outage - BLRM recovery AA.jpgADC.jpg
                      Reply  Tue Apr 17 09:57:09 2018, Steve, Update, General, seismometer interfaces BS_Tril_Intrf-1X5.jpgGurs_Intf-1X1.jpg
                         Reply  Wed Apr 18 20:33:19 2018, Kevin, Update, General, seismometer interfaces front.jpgback.jpgconnectors.jpg
                            Reply  Fri Apr 20 23:36:28 2018, Kevin, Update, PEM, Seismometer BLRMs front.jpgback.jpgdetail.jpg
                               Reply  Tue Apr 24 21:19:08 2018, Kevin, Update, PEM, Seismometer BLRMs modified_schematic.pdftop.jpgbottom.jpgBS_Seis_PSD.pdf
                                  Reply  Wed Apr 25 17:44:39 2018, Arnold, Update, PEM, PEM Anti-Alias wiring 
                                     Reply  Thu Apr 26 09:35:49 2018, Kevin, Update, PEM, PEM Anti-Alias wiring board.jpgSeismometerPSD.pdf
                                        Reply  Thu Apr 26 20:22:21 2018, Kevin, Update, PEM, ADC common mode rejection with new seismometer connections setup.pdfseismometerASD.pdf
                                           Reply  Wed May 16 21:02:22 2018, Kevin, Update, PEM, ADC common mode rejection with new seismometer connections ADC_noise.pdf
                                        Reply  Thu Jun 14 15:24:32 2018, Steve, Update, PEM, ADC DAC In Line Test Boards are in ADC_DAC_in_(1).JPG
    Reply  Tue Jan 2 15:43:35 2018, Steve, Update, VAC, pumpdown after power outage pumpdown_from_44_mTorr.png
       Reply  Wed Feb 7 15:50:42 2018, Steve, Update, VAC, IFO pressure monitoring CC_analog.png
Message ID: 13510     Entry time: Sat Jan 6 18:27:37 2018     In reply to: 13507     Reply to this: 13755
Author: gautam 
Type: Update 
Category: General 
Subject: power outage - IFO recovery 

Mostly back to nominal operating conditions now.

  1. EX TransMon QPD is not giving any sensible output. Seems like only one quadrant is problematic, see Attachment #1. I blame team EX_Acromag for bumping some cabling somewhere. In any case, I've disabled output of the QPD, and forced the LSC servo to always use the Thorlabs "High Gain" PD for now. Dither alignment servo for X arm does not work so well with this configuration - to be investigated.
  2. BS Seismometer (Trillium) is still not giving any sensible output.
    • I looked under the can, the little spirit level on the seismometer is well centered.
    • I jiggled all the cabling to rule out any obvious loose connections - found none at the seismometer, or at the interface unit (labelled D1002694 on the front panel) in 1X5/1X6.
    • All 3 axes are giving outputs with DC values of a few hundred - I guess there could've been some big earthquake in early December which screwed the internal alignment of the sensing mass in the seismometer. I don't know how to fix this.
    • Attachment #2 = spectra for the 3 channels. Can't say they look very seismicy frown. I've assumed the units are in um/sec.
    • This is mainly bothering me in the short term because I can't use the angular feedforward on PRC alignment, which is usually quite helpful in DRMI locking.
    • But I think the PRM Oplev loop is actually poorly tuned, in which case perhaps the feedforward won't really be necessary once I touch that up.

What I did today (may have missed some minor stuff but I think this is all of it):

  1. At EX:
    • Toggled power to Thorlabs trans monitoring PD, checked that it was actually powered, squished some cables in the e- rack.
    • Removed PDA55 in the green path (put there for EX laser AM/PM measurement). So green beam can now enter the X arm cavity.
    • Re-connected ALS cabling.
    • Turned on HV supply for EX Green PZT steering mirrors (this has to be done every time there is a power failure).
  2. At ITMY table:
    • Removed temporary HeNe RIN/ Oplev sensing noise measurement setup. HeNe + 1" vis-coated steering mirror moved to SP table.
    • Turned on ITMY/SRM Oplev HeNe.
    • Undid changes on ITMY Oplev QPD and returned it to its original position.
    • Centered ITMY reflected beam on this QPD.
  3. At vertex area
    • Looked under Trillium seismometer can - I've left the clamps undone for now while we debug this problem.
    • Power-cycled Trillium interface box.
    • Touched up PMC alignment.
  4. Control room
    • Recover IFO alignment using combination of IR and Green beams.
    • Single arm locking recovered, dither alignment servos run to maximize arm transmission. Single arm locks holding for hours, that's good.
    • The X arm dither alignment isn't working so well, the transmission never quite hits 1 and it undergoes some low frequency (T~30secs) oscillations once the transmission reaches its peak value.
    • Had to do the usual ipcrm thing to get dataviewer to run on pianosa.

Next order of business:

  1. Recover ALS:
    • aim is to replace the vertex area ALS signals derived from 532nm with their 1064nm counterparts.
    • Need to touch up end PDH servos, alignment/MM into arms, and into Fibers at ends etc.
    • Control the arms (with RMs misaligned) in the CARM/DARM basis using the revised ALS setup.
    • Make a noise budget - specifically, we are interested in how much actuation range is required to maintain DARM control in this config.
  2. Recover DRMI locking
    • Continue NBing.
    • Do a statistical study of actuation range required for acquiring and maintaining DRMI locking.
Attachment 1: EX_QPD_Quad1_Faulty.pdf  69 kB  | Hide | Hide all
EX_QPD_Quad1_Faulty.pdf
Attachment 2: Trillium_faulty.pdf  37 kB  | Hide | Hide all
Trillium_faulty.pdf
ELOG V3.1.3-